OpenStudy (anonymous):

What is the volume of the coffee in a can if the can has a radius of 4 inches and a height of 9 inches? (Use 3.14 for π.)

OpenStudy (anonymous):

113.04 in.3 226.08 in.3 339.12 in.3 452.16 in.3

OpenStudy (anonymous):

226.08 inches

OpenStudy (mathstudent55):

The volume of a cylinder is given by the formula: V = (pi)(r^2)h, where V = volume, pi = 3.14, r = radius, h = height In this question, V = 3.14(4^2)(9) = 452.16 in^3